Determine the frequency of sound heard by a stationary listener when a moving siren emits 550Hz and travels away at 50m/s from the listener assuming the air temperature is 20 degree Celsius.




1
Expert's answer
2022-04-20T10:13:02-0400
"f=f_0\frac{c}{c+v}"

"f=550*\frac{343}{343+50}=480\;\rm Hz"